# A universal scheme to self-test any quantum state and measurement

### v1

@@ A.1 Bell inequalities
where to arrive at the last line we used the fact that $\tilde{A}_{i,j}^{2}\leqslant\mathbbm{1}$. Thus, one can conclude from (26) that $3(N-1)$ is an upper bound on $\mathcal{I}_{l}$ when using quantum theory. One can easily verify that the Bell inequalities (17) attain the value $3(N-1)$ when the states shared among the parties are the GHZ-like state $|\phi_{l}\rangle$ (20) and the observables given in Eq. (19). Thus, the Tsirelson’s bound of the Bell inequalities (17) for any $l$ is $3(N-1)$. ∎
@@
Also, one can conclude from Eq. (26) that the observables that attain the maximum violation of the Bell inequalities (17) must be unitary, that is, $A_{i,j}^{2}=\mathbbm{1}$. These relations will be particularly useful for self-testing.
@@ D.  Measurements $A_{i,2}$
Appendix B Appendix B: Self-testing any projective measurement
@@
In this section, we show how to certify Eve’s second measurement $E_{1}=\{R_{l|1}\}$ to be a particular reference projective measurement $E^{\prime}_{1}=\{R^{\prime}_{l|1}\}$ (up to some additional degrees of freedom). For this purpose, we use the states generated by the sources, certified as in Eq. (30), and the certified measurements in Eqs. (31) and (32) along with some additional conditions imposed on the observed correlations.
@@
Let us then decompose every measurement element in the Pauli basis,
as
@@
[…]
@@
Appendix C Appendix C: Self-testing any rank-one extremal quantum measurement
In this section, we show that one can self-test the measurement corresponding to the input $e=1$ with the central party to be a general extremal rank-one quantum measurement $E_{1}=\{R^{\prime}_{l|1}\}$ up to some additional degrees of freedom. For this purpose, we use the states generated by the sources $P_{i}$, certified as in Eq. (29), and the certified measurements in Eqs. (31) and (32) along with some additional conditions.
Let us then consider an ideal reference measurement $E^{\prime}_{1}=\{R^{\prime}_{l|1}\}$ with $R^{\prime}_{l|1}$ $(l=1,\ldots,K)$ being measurement elements defined on some Hilbert space $\mathbbm{C}^{D}$ and assume this measurement to be extremal. For completeness let us recall that an extremal POVM is one that cannot be decomposed into a convex combination of other POVM’s. Additionally, we assume that the measurement elements are rank-one, in which case they can be expressed as
$R^{\prime}_{l|1}=\lambda_{l}|\delta_{l}\rangle\!\langle\delta_{l}|,$ (138)
where $0\leqslant\lambda_{l}\leqslant 1$ for any $l$ with the property that the states $\{|\delta_{l}\rangle\}$ are linearly independent [cf. Ref. [34]].
Let us now notice that similarly to the case of projective measurements discussed above, the extremal rank-one POVM $\{R^{\prime}_{l|1}\}$ can also be embedded in $(\mathbbm{C}^{2})^{\otimes N}$. To this end we consider the minimal $N$ such that $D\leq 2^{N}$ and associate to every element of the standard basis $|i\rangle\in\mathbbm{C}^{D}$ a vector $|i_{1}\ldots i_{N}\rangle$ forming the standard basis in $\in(\mathbbm{C}^{2})^{\otimes N}$ such that $i_{1}\ldots i_{N}$ is the binary representation of $i=0,\ldots,D-1$. Since then the measurement elements do not sum up to the identity on $(\mathbbm{C}^{2})^{\otimes N}$, we need to consider additional measurement elements $J_{k}$ $(k=1,\ldots,2^{N}-D)$ such that
$\sum_{l=1}^{K}R^{\prime}_{l|1}+\sum_{k=1}^{2^{N}-D}J_{k}=\mathbbm{1}_{2^{N}}$ (139)
We take this additional operators to be rank-one projections, $J_{k}=|\varphi_{k}\rangle\!\langle\varphi_{k}|$ where $|\varphi_{k}\rangle$ are mutually orthogonal. Moreover, the vectors $|\varphi\rangle$ are orthogonal to $\sum_{l=1}^{K}R^{\prime}_{l|1}$. Such vectors can always be found, for instance by taking an arbitrary orthonormal basis in the subspace of $(\mathbbm{C}^{2})^{\otimes}$ which is orthogonal to the support of $\sum_{l=1}^{K}R^{\prime}_{l|1}$. Let us denote the new measurement by $N=\{\lambda_{1}|\delta_{1}\rangle\!\langle\delta_{1}|,\ldots,\lambda_{K}|\delta_{K}\rangle\!\langle\delta_{K}|,J_{1},\ldots,J_{2^{N}-D}\}$ and notice that it is extremal too because it is obtained by complementing the extremal measurement $E_{1}^{\prime}$ by rank-one projections [34].
Let us now represent any measurement element $N_{l}$ in terms of the Pauli basis (108) as
Notice that the above statistics can be realised if the sources generate the two-qubit maximally entangled state $|\phi^{+}\rangle=(1/\sqrt{2})(|00\rangle+|11\rangle)$ and the measurement $E_{1}=\{R_{l|1}\}$ is exactly $R_{l|1}=(R^{\prime}_{l|1})^{*}$ where $R^{\prime}_{l|1}$ is given in (138) and *** denotes the complex conjugate, whereas the external parties perform the measurements in Eq. (19).
Theorem 3.
[…]
Proof.
The rest of the proof follows exactly from Eqs. E13-E29 of [35]. One first needs to decompose the state $\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}$ in its eigenbasis, let’s say $\{|k\rangle\}$, and then use the fact that extremal POVM’s can not be decomposed into other POVM’s to obtain that
$\langle k|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|k\rangle=f^{l}_{i_{1},\ldots,i_{N}}\qquad\forall k,i_{1},\ldots,i_{N}.$ (149)
Then, using the fact that the measurement elements of extremal POVM’s are linearly independent, one can conclude that for all $i_{1},\ldots,i_{N}$
$\langle s|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|k\rangle=0\qquad\forall s,k\quad\mathrm{s.t.}\quad s\neq k$ (150)
which allows us to conclude that over the support of the state $\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}$
$\tilde{R}^{l}_{i_{1},\ldots,i_{N}}=f^{l}_{i_{1},\ldots,i_{N}}\mathbbm{1}_{E^{\prime\prime}}$ (151)
Similarly, when $A_{i,2}=-Y\otimes\mathbbm{1}$ for all $i$, then we will obtain exactly the similar relation to Eq. (148) as
$\mathrm{Tr}\left[\left(\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}\right)\tilde{R}_{l,i_{1},\ldots,i_{N}}\right]=g(i_{1},\ldots,i_{N})f_{l,i_{1},\ldots,i_{N}}\quad\forall i_{1},\ldots,i_{N}$ (152)
where $g(i_{1},\ldots,i_{N})=1$ for even number of $j$ and $g(i_{1},\ldots,i_{N})=-1$ for odd number of $j$ such that $i_{j}=2$. Thus, we can certify any extremal POVM. ∎
Appendix D Appendix D: Self-testing any quantum state
Using the Theorems 2 and 3 proven above, we can now show how it is possible to probabilistically self-test arbitrary pure or mixed $N$-qubit quantum state. For this purpose, let us first certify the post-measurement state when eve chooses $e=1$ and obtains an outcome $l$.
@@ Appendix D Appendix D: Self-testing any quantum state
Assume that the states are certified as in Eq. (30) and Eve’s measurement $E_{1}$ is certified as in Eqs. (2) and (3). Consequently, when Eve observes the $l-$th outcome of her measurement, the post-measurement state with the external parties is given by
[…]

### v2

@@ A.1 Bell inequalities
where to arrive at the last line we used the fact that $\tilde{A}_{i,j}^{2}\leqslant\mathbbm{1}$. Thus, one can conclude from (A.1) that $3(N-1)$ is an upper bound on $\mathcal{I}_{l}$ when using quantum theory. One can easily verify that the Bell inequalities (13) attain the value $3(N-1)$ when the states shared among the parties are the GHZ-like state $|\phi_{l}\rangle$ (16) and the observables given in Eq. (15). Thus, the Tsirelson’s bound of the Bell inequalities (13) for any $l$ is $3(N-1)$. ∎
@@
Also, one can conclude from Eq. (A.1) that the observables that attain the maximum violation of the Bell inequalities (13) must be unitary, that is, $A_{i,j}^{2}=\mathbbm{1}$. These relations will be particularly useful for self-testing.
@@ D.  Measurements $A_{i,2}$
Appendix B Appendix B: Self-testing any extremal POVM
Let us then consider an ideal reference measurement $E^{\prime}_{1}=\{R^{\prime}_{l|1}\}$ with $R^{\prime}_{l|1}$ $(l=1,\ldots,K)$ being measurement elements defined on some Hilbert space $\mathbbm{C}^{D}$ and assume this measurement to be extremal. For completeness let us recall that an extremal POVM is one that cannot be decomposed into a convex combination of other POVM’s. As $R^{\prime}_{l|1}$ are of arbitrary rank, let denote them by $r_{l}=\mathrm{rank}(R^{\prime}_{l|1})$. Let us also assume that the measurement has $K\leq 2^{N}$ outcomes; $r_{1}+\ldots+r_{K}=D$. Such a measurement can always be embedded in an $N$-qubit Hilbert space $(\mathbbm{C}^{2})^{\otimes N}$ where $N$ the minimal natural number such that $D\leq 2^{N}$. In order to represent $R^{\prime}_{l|1}$ in the $N$-qubit Hilbert space one can for instance assign to any element of the standard basis $|i\rangle\in\mathbbm{C}^{D}$ an element $|i_{1}\ldots i_{N}\rangle$ of the product standard basis of $(\mathbbm{C}^{2})^{\otimes N}$ such that $i_{1}\ldots i_{N}$ is a binary representation of $i$ $(i=0,\ldots,D-1)$. Then, one needs to add another projective measurement element $M^{\perp}$ to $E_{1}^{\prime}$ which is orthogonal to the support of $E_{1}^{\prime}$ so that the resulting measurement defined on $(\mathbbm{C}^{2})^{\otimes N}$ satisfies
$\sum_{i=1}^{K}R^{\prime}_{i|1}+M^{\perp}=\mathbbm{1}_{2^{N}},$ (100)
where $\mathbbm{1}_{2^{N}}$ is the identity acting on $(\mathbbm{C}^{2})^{\otimes N}$ where $\mathrm{rank}(M^{\perp})=2^{N}-D$. Let us denote the new $(K+1)-$outcome measurement by $N=\{R^{\prime}_{1|1},\ldots,R^{\prime}_{K|1},M^{\perp}\}$ and notice that it is extremal too because it is obtained by complementing the extremal measurement $E_{1}^{\prime}$ by a projector orthogonal to its support [34].
Let us now represent any measurement element $N_{l}$ in terms of the Pauli basis
as
Notice that the above statistics can be realised if the sources generate the two-qubit maximally entangled state $|\phi^{+}\rangle=(1/\sqrt{2})(|00\rangle+|11\rangle)$ and the measurement $E_{1}=\{R_{l|1}\}$ is exactly $R_{l|1}=(R^{\prime}_{l|1})^{*}$ where $R^{\prime}_{l|1}$ are the reference measurement operators and * denotes the complex conjugate, whereas the external parties perform the measurements in Eq. (15).
Theorem 2.
$U_{E}R_{l|1}U_{E}^{\dagger}$ $=$ $(R^{\prime}_{l|1})^{*}\otimes\mathbbm{1}_{E^{\prime\prime}},\quad\mathrm{if}\quad A_{i,2}=Y\otimes\mathbbm{1}\quad\forall i,\quad\text{or}\quad$
Proof.
The rest of the proof follows the same lines as the proof of Theorem 6 of Ref. [35]. We first decompose the state $\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}$ in its eigenbasis, let’s say $\{|b\rangle\}$ with eigenvalues $p_{b}$ which allows us to obtain from Eq. (110) that
$\sum_{b}p_{b}\langle b|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|b\rangle=f^{l}_{i_{1},\ldots,i_{N}}.$ (111)
Next, we introduce a collection of POVMs $\{R_{l}^{b}\}_{l}$ for any $b$, the measurement operators of which are given by
$R_{l}^{b}$ $=$ $\mathrm{Tr}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\left[\left(\mathbbm{1}_{A^{\prime}}\otimes|b\rangle\!\langle b|_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\right)\overline{R}_{l}^{T}\right]$ (112)
$=$ $\sum_{\textbf{}i_{1},\ldots,i_{N}=0}^{3}\langle b|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|b\rangle\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}.$
Since $\overline{R}_{l}^{T}\geq 0$ it directly follows from the above equation that $R_{l}^{b}\geq 0$ for any $l$ and $b$. Moreover, $\sum_{l}\overline{R}_{l}^{T}=\mathbbm{1}$ implies that $\sum_{l}R_{l}^{b}=\mathbbm{1}$ for any $b$. Thus, $\{R_{l}^{b}\}_{l}$ are proper quantum measurements for all $b$.
These additional POVMs, through Eq. (111), allow us to decompose the ideal POVM $\{N_{l}\}$ as $N_{l}=\sum_{b}p_{b}R_{l}^{b}$. As the ideal POVM is extremal, it can not be decomposed in terms of other POVM’s. Consequently, we have that
$\forall b\qquad R_{l}^{b}=N_{l},$ (113)
which is equivalent to
$\langle b|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|b\rangle=f^{l}_{i_{1},\ldots,i_{N}}\qquad\forall b.$ (114)
Let us now consider the following vectors:
$|\varphi_{a,s,t}\rangle=\frac{1}{\sqrt{2}}\left(|s\rangle\pm\mathbbm{i}^{a}|t\rangle\right),$ (115)
such that $|s\rangle$ and $|t\rangle$ are vectors that belong to the eigenbasis $\{|k\rangle\}$ of $\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}$ such that $s\neq t$ and $a=0,1$. Let us look at the quantity
$\mathrm{Tr}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\left[\left(\mathbbm{1}_{A^{\prime}}\otimes|\varphi_{a,s,t}\rangle\!\langle\varphi_{a,s,t}|_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\right)\overline{R}_{l}^{T}\right]=\sum_{i_{1},\ldots,i_{N}=0}^{3}\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\ \mathrm{Tr}(|\varphi_{a,s,t}\rangle\!\langle\varphi_{a,s,t}|_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\tilde{R}^{l}_{i_{1},\ldots,i_{N}}),$ (116)
which with the aid of the explicit form of the vector (115) can be rewritten as
$\mathrm{Tr}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\left[\left(\mathbbm{1}_{A^{\prime}}\otimes|\varphi_{a,s,t}\rangle\!\langle\varphi_{a,s,t}|_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\right)\overline{R}_{l}^{T}\right]=N_{l}\pm\mathrm{Tr}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\left[(\mathbbm{1}_{A^{\prime}}\otimes L^{a}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}})\overline{R}_{l}^{T}\right],$ (117)
where $L_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}^{a}=(\mathbbm{i}^{a}/2)\left(|t\rangle\!\langle s|+(-1)^{a}|s\rangle\!\langle t|\right)$. Using the fact that $\overline{R}_{l}^{T}$ is positive semi-definite, which implies that the left-hand side of the above relation is positive semi-definite too, we conclude that
$N_{l}\geq\pm\mathrm{Tr}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\left[(\mathbbm{1}_{B^{\prime}}\otimes L^{a}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}})\overline{R}_{l}^{T}\right]=:\pm\Omega_{l}^{a}.$ (118)
As we show at the end of the proof, the above operator inequality can be satisfied only if $\mathrm{supp}(\Omega_{l}^{a})\subseteq\mathrm{supp}(N_{l})$ for any $l$ and for both values of $a=0,1$. Let us then denote by $|\phi_{l,m}\rangle$ $(m=1,\ldots,r_{l})$ as the eigenvectors of $N_{l}$, where $r_{l}$ denotes the rank of $N_{l}$. Due to the fact that $\mathrm{supp}(\Omega_{l}^{a})\subseteq\mathrm{supp}(N_{l})$, each operator $\Omega_{l}^{a}$ can be written using the eigenvectors of $N_{l}$ as
$\Omega_{l}^{a}=\sum_{m,n=1}^{r_{l}}\gamma_{m,n}^{(l,a)}|\phi_{l,m}\rangle\!\langle\phi_{l,n}|,$ (119)
where $\gamma_{m,n}^{(l)}$ are some complex coefficients; recall that $\Omega_{l}^{a}$ are Hermitian.
Let us now notice that it follows from Eq. (118) that
$\sum_{l}\Omega_{l}^{a}$ $=$ $\sum_{l}\mathrm{Tr}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\left[(\mathbbm{1}_{A^{\prime}}\otimes L^{a}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}})\overline{R}_{l}^{T}\right]$ (120)
$=$ $\mathrm{Tr}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\left[(\mathbbm{1}_{A^{\prime}}\otimes L^{a}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}})\right]=0,$
where the second equality follows from the fact that $\overline{R}_{l}^{T}$ sum up to the identity, whereas the third equality is a consequence of the fact that the $L^{a}$ operators are traceless. We thus have that for any $a$,
$\sum_{l}\Omega_{l}^{a}=0.$ (121)
After plugging Eq. (119), we obtain the following equation
$\sum_{l,m,n}\gamma_{m,n}^{(l,a)}|\phi_{l,m}\rangle\!\langle\phi_{l,n}|=0$ (122)
Let us now exploit the fact that the measurement $\{N_{l}\}$ is extremal. As proven in Ref. [34] it follows from the latter that the operators $|\phi_{l,m}\rangle\!\langle\phi_{l,n}|$ are linearly independent for all $l$, $m$ and $n$. It thus follows from Eq. (122) that $\gamma$’s vanish and therefore
$\Omega_{l}^{a}=\mathrm{Tr}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\left[(\mathbbm{1}_{A^{\prime}}\otimes L^{a}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}})\overline{R}_{l}^{T}\right]=0$ (123)
for any $a$ and any $l$. Using now the explicit form of the operators $L^{a}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}$, the above implies the following conditions for the $\overline{R}_{l}^{T}$ operators,
$\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\left(\langle s|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|t\rangle+\langle t|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|s\rangle\right)=0,$ (124)
for $a=0$, and
$\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\left(\langle t|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|s\rangle-\langle s|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|t\rangle\right)=0$ (125)
for $a=1$, where we used the fact that $\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}$ are linearly independent for all $i_{1},\ldots,i_{N}$. The only possible solution to these equations is that
$\langle s|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|t\rangle=0\qquad(s\neq t).$ (126)
Combining this with Eq. (114) we obtain that
$\tilde{R}^{l}_{i_{1},\ldots,i_{N}}=f_{i_{1},\ldots,i_{N}}^{l}\mathbbm{1},$ (127)
which implies the desired relation $\overline{R}_{l}^{T}=N_{l}\otimes\mathbbm{1}_{A_{1}^{\prime\prime}\ldots A_{N}^{\prime\prime}}$ for all $l$.
Let us finally prove that $\mathrm{supp}(\Omega_{l}^{a})\subseteq\mathrm{supp}(N_{l})$ for any $l$ and $a$. For this purpose, let us pick a particular $N_{l}$ and consider its eigenvectors $|\psi_{i}\rangle$ $(i=1,\ldots,r_{l})$, where $r_{l}$ is the rank of $N_{l}$. We then construct an orthonormal basis $\{|\phi_{i}\rangle\}_{i}$ with $i=1,\ldots,2^{N}$ in the $N$-qubit Hilbert space such that $|\phi_{i}\rangle=|\psi_{i}\rangle$ for $i=1,\ldots,r_{l}$.
Sandwiching then inequalities (118) with $|\phi_{i}\rangle$ and $|\phi_{j}\rangle$, one obtains
$\forall_{i,j}\qquad\langle\phi_{i}|N_{l}|\phi_{j}\rangle\geq\pm\langle\phi_{i}|\Omega_{l}^{a}|\phi_{j}\rangle.$ (128)
Now, for any pair $i,j$ such that $i\geq r_{l}+1$ or $j\geq r_{l}+1$ (i.e., either $|\phi_{i}\rangle$ or $|\phi_{j}\rangle$ belongs to the kernel of $N_{l}$), the above condition gives us
$0\geq\pm\langle\phi_{i}|\Omega_{l}^{a}|\phi_{j}\rangle,$ (129)
which directly implies that $\langle\phi_{i}|\Omega_{l}^{a}|\phi_{j}\rangle=0$ for any pair $i,j$ such that either $i\geq r_{l}+1$ or $j\geq r_{l}+1$. This directly implies that $\Omega_{l}^{a}$ act nontrivially only on $\mathrm{supp}(N_{l})$, and thus $\mathrm{supp}(\Omega_{l}^{a})\subseteq\mathrm{supp}(N_{l})$.
B.1 Alternative proof for projective measurements
@@
Here we present an alternative proof in the case when we want to certify Eve’s second measurement $E_{1}=\{R_{l|1}\}$ to be a particular reference projective measurement $E^{\prime}_{1}=\{R^{\prime}_{l|1}\}$ (up to some additional degrees of freedom). For this purpose, we use the states generated by the sources, certified as in Eq. (23), and the certified measurements in Eqs. (24) and (25) along with some additional conditions imposed on the observed correlations.
@@
Let us then decompose every measurement element in the Pauli basis (101), as
@@
$U_{E}R_{l|1}\,U_{E}^{\dagger}$ $=$ $(R^{\prime}_{l|1})^{*}\otimes\mathbbm{1}_{E^{\prime\prime}},\quad\mathrm{if}\quad A_{i,2}=Y\otimes\mathbbm{1}\quad\forall i,\quad\text{or}\quad$
@@
Appendix C Appendix C: Self-testing any quantum state
Using the Theorems 2 proven above, we can now show how it is possible to probabilistically self-test arbitrary pure or mixed $N$-qubit quantum state. For this purpose, let us first certify the post-measurement state when eve chooses $e=1$ and obtains an outcome $l$.
@@ Appendix C Appendix C: Self-testing any quantum state
Assume that the states are certified as in Eq. (23) and Eve’s measurement $E_{1}$ is certified as in Eq. (2). Consequently, when Eve observes the $l-$th outcome of her measurement, the post-measurement state with the external parties is given by
$U_{A}\,\rho_{A}^{l}\,U_{A}^{\dagger}$ $=$ $\frac{1}{\mathrm{Tr}R^{\prime}_{l|1}}R^{\prime}_{l|1}\otimes\tilde{\rho}_{A^{\prime\prime}},\quad\mathrm{if}\quad A_{i,2}=Y\otimes\mathbbm{1}\quad\forall i,\quad\text{or}\quad$